OpenStudy (anonymous):

Find a so that the point (-1,2) is on the graph of f(x)=ax^2+5

jimthompson5910 (jim_thompson5910):

(x,y) = (-1,2) so x = -1 and y = 2

jimthompson5910 (jim_thompson5910):

f(x)=ax^2+5 y = a*x^2 + 5 2=a*(-1)^2+5 ... plug in x = -1 and y = 2 2 = a*1 + 5 solve for 'a'

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

yep